Ticket Number: HEA-NOC/20080527-9 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: e-net
Ticket Opened: 20080527 14:44 UTC+1 Problem Start: 20080528 19:00 UTC+1
Ticket Update: 20080529 10:00 UTC+1 Problem End: 20080529 06:00 UTC+1
Site/line: CAO, GMIT, IT Sligo, LYIT, NUIG Clinical Sciences
e|net will be carrying out fibre splicing work in Galway on Wednesday 28th
May into Thursday 29th May. This is to splice in two new NBE connections.
Loss of resilience to the Galway ring for two one hour periods during the
eleven hour window. No actual outages are expected.
20080527 14:36 BN Confirmed with e|net that the work was ok to go ahead on
such short notice as no outages will occur.
20080529 01:42 COC E-net have started work on this at 01:20
20080529 03:42 COC Fibre ring has been closed again as of 03:18. One
possible problem between 2nd Clinical Science link. E-net are
20080529 09:23 BN This work has been completed, however there are some
outstanding issues. One leg of the Clinical Sciences connection is still
down due to excess attenuation. This should be resolved by NUIG this
morning. During the work there appeared to be extremely brief (< 1min)
outages to IT Sligo and LYIT. These outages were not expected and the
causes are being investigated.
20080529 09:59 BN The attenuators in the Clinical Sciences site have been
removed and both connections are now working as expected.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080527-9
HEAnet Limited. Registered in Ireland, No. 275301.